OpenStudy (anonymous):

find the area inbetween x^2+6 and 12-x^2

OpenStudy (zepp):

We could scale these two functions down so the intersection points lie on the x-axis.

OpenStudy (zepp):

Since I'm quite lazy, let's just put them in Wolfram|Alpha and http://puu.sh/1qVS8 Intersection points are \((\sqrt3, 9)\) and \((-\sqrt3, 9)\)

OpenStudy (zepp):

So let's move these two functions downwards, and we get y=x^2-3 and y=3-x^2

OpenStudy (zepp):

Then since our two functions look like this, we can take the integral of both functions from -sqrt3 to sqrt3 and add them up

OpenStudy (anonymous):

thx i got that one..im on another now that i used wolfram aplia and when i plug it into my homework website it marks it as wrong...i have 1 chance left

OpenStudy (zepp):

\[\large \int_{-\sqrt{3}}^{\sqrt{3}}x^2-3dx+\int_{-\sqrt{3}}^{\sqrt{3}}3-x^2dx\]

OpenStudy (anonymous):

its the area between 1/x and 1/(x^2) between .5 and 16

OpenStudy (zepp):

Umm, let's ask @TuringTest for this one :s

OpenStudy (anonymous):

if im correct they integrate @ 1 so i put in integrate (1/(x^2) - (1/x) from .5 to 1 and got .0306853 integrate (1/x) - (1/(x^2) from 1 to 16 and got log 16 - (15/6) i added them together for the approximate answer i put 1.8657740222 and for the exact i put (ln 16 - (15/6)) + .0306853

OpenStudy (anonymous):

do i go to his page and ask him?

OpenStudy (zepp):

He's probably busy atm, hold on a second

OpenStudy (zepp):

Sorry, you'll have to wait a little bit, I don't know how to do that one :/

OpenStudy (anonymous):

its ok.....but im waiting for you to talk to him right?

OpenStudy (zepp):

Split the atom.

OpenStudy (zepp):

@AccessDenied will answer your question, while he's typing out the question, lemme post a plot of this situation: In blue: 1/x, in red: 1/x^2

OpenStudy (accessdenied):

I think your approach is correct (I was originally going to type that explanation), although I seem to get a different value in calculator than you specified for the first integral. I get about the same answer, except one decimal place to the right (0.306...).